The square root of 8y times the square root of 9y

Answer by Alan3354(69443)

The square root of 8y times the square root of 9y
========
--> sqrt(8y*9y) = sqrt(72y^2)
= sqrt(36y^2)*sqrt(2)
= 6y*sqrt(2)